OpenStudy (anonymous):

Which of the following points are NOT on the graph of y = x + 2?

OpenStudy (anonymous):

HELP ME

OpenStudy (anonymous):

Probably need the list of points. For each ordered pair, put the x and y into the equation. If it results in a true statement, the point is on the line. If not, not.

OpenStudy (phi):

if you have a (x,y) pair of numbers, replace the x and y in the equation and see if it is true.

OpenStudy (phi):

It is arithmetic.

OpenStudy (phi):

for example if you have (0,2) replace x with 0 and y with 2 in y= x+2 2= 0+2 2=2 so that pair worked.

OpenStudy (phi):

can you do the problem?